### v3

@@ Abstract
Index Terms:
Complex-Valued Neural Networks, Kolmogorov-Arnold Networks, Explainable AII Introduction
The recently published Kolmogorov-Arnold Network (KAN) [1, 2] has proven to be a successful new approach for problems especially in the field of symbolic function fitting [3] and solving physical equations [4]. Adding to its success is the intrinsic explainability of KANs because of learnable univariate functions on edges instead of linear weights and fixed nonlinear activation functions in classical Multilayer Perceptrons. These learned univariate functions hold all model weights and can easily be visualized, thus making it intuitively understandable how single layers impact the model output.
Another emerging field in machine learning research are Complex-Valued Neural Networks (CVNNs), which have shown great success in a multitude of applications as well as good theoretical properties in terms of overfitting and generalization [5, 6]. Especially for applications with complex-valued input types, these models have shown promising results.
@@
The original KANs used fitting of B-Splines to learn the edge functions, however this training is rather slow and can be unstable, thus [7] proposed to replace the originally used B-Splines with easier to learn Radial Basis Functions. While it would be possible to utilize complex-valued B-Splines [8, 9, 10] to construct a complex-valued KAN, we propose to adopt the idea of RBFs to the complex domain, to benefit from their easier and faster training process. Additionally, we propose to replace the tedious dynamic grid extension used in KANs by Batch Normalization to prevent the model to fall outside the predefined grid.
@@ II-A Kolmogorov-Arnold Networks
After the recent introduction of KANs by Liu et al. [1, 2] there has been a great variety of attempts to apply the ideas of KANs to different fields like image processing [11], satellite image segmentation [12], graph neural networks [13, 14] and even transformers [15]. Hou et al. [16] give a detailed overview of the different applications and extensions of KANs. Their explainability can be of great value in fields where machine learning approaches are strongly regulated like survival analysis in medicine or engineering [17]. Multiple works have already benchmarked the performance of KANs against MLPs [3, 18] and found KANs to be a more suitable alternative in some fields. Alter et al. [19] found that large-scale KANs are more robust against adversarial attacks as MLPs and thus form an interesting direction for further research in multiple fields.
In [4] the authors explore different partial differential equation forms based on KAN instead of MLP for solving forward and inverse problems in computational physics. A systematical comparison demonstrates that the KAN approach significantly outperforms MLP regarding accuracy and convergence speed. Further successful applications of KAN can be found for operator learning in computational mechanics [20] and image classification [21].
@@ II-B Complex-Valued Neural Networks
After early introduction of CVNNs [22, 23, 24, 25] they have lately risen in popularity since the introduction of building blocks for deep learning architectures [26]. Since then a lot of theoretical contributions have been made [27, 28, 29] to enable a multitude of applications [30, 31, 32, 33, 34].
The most closely related prior work in the complex domain are deep Complex Radial Basis Function Neural Networks [35, 36]. However, in CRBFNs, the RBFs are applied to all inputs of a neuron (e.g. vertex of the computational graph) simultaneously, while in CVKANs the RBFs are applied on the edges of the computational graph to each value individually. Thus CRBFNs are architecturally more similar to classical MLPs with RBFs as activation functions, where we instead aim to adopt the KAN framework to the complex domain.
@@ III Real-Valued Kolmogorov-Arnold Networks
The Kolmogorov-Arnold representation theorem (1) [37] states that any continuous function can be represented by a superposition of univariate functions.
This formula can be used to construct a two layer neural network with inputs, one output and learnable functions . For each layer and each feature an individual univariate function has to be learned, so combination of features is only done by summing the outputs of all these functions. While theoretically possible, this theorem was long deemed to be of little use for machine learning [38], because the inner functions can be highly nonsmooth.
However, Liu et al. [1] have generalized this formulation of only one hidden layer of size and one single output to arbitrary network depths and layer widths with to overcome the problem of nonsmooth inner functions. Let be the -th neuron in layer , then every node in layer is connected to every node in the following layer by an edge with and . On every edge there is one learnable univariate function and the values of all incoming edges into a single node of the next layer are summed. This way the multiplication with linear weights between layers in classical MLPs has been replaced by learnable univariate functions and the nonlinear but fixed activation functions in the nodes of MLPs have been replaced by plain summation. In the original KAN the functions are learned using B-Splines [1, 2].
@@ III-A Real-Valued Radial Basis Functions
The way B-Splines are used to learn the functions in KANs [1, 2] is computationally intensive and slow. Thus different improved methods for learning a function have been proposed like DeepOKAN [20] or FastKAN [7], in which the authors suggest to replace B-Splines with RBFs for faster computation. Equation (2) describes a RBF that is symmetrical and centered around .
@@ IV-A Complex-Valued Radial Basis Functions
Chen et al. [35] have chosen to treat real and imaginary parts separately and construct a complex-valued RBF out of two real-valued RBFs which only depend on the real or imaginary part of the input each. On the other hand Soares et al. [36] use a single real-valued RBF that depends on both the real and imaginary parts of the input jointly. While the latter apply these multivariate RBFs inside vertices, we propose to use univariate RBFs on edges instead to maintain the idea of KANs and achieve higher explainability. We define a as:
@@ IV-B Residual Activation Function
We propose to use SiLU (8) as this residual activation function, a complex-valued equivalent to the SiLU-function (7) used in [1, 2, 7]:
@@ IV-C Complex-Valued Batch Normalization
For KANs it is necessary to fix the grid range. However, it is not guaranteed that the output of an intermediate layer (and thus the input of the next layer) stays within a grid that was fixed before training. To tackle this challenge [1] proposed dynamic extension of the grid during training. This is however a time intensive process that might slow down the learning process. Instead we propose to normalize the output of a vertex after summation. We propose to use BatchNorm () [39] for the Radial Basis Function based FastKAN [7]. To adapt this to the complex domain, we explore three different complex-valued Batch Normalization approaches.
First, we adapt BatchNorm [26], where the covariance matrix of the complex distribution is normalized and an output distribution is learned through a learnable covariance (e.g. a symmetric positive definite) matrix. With being the mean and covariance over the batch dimension and learnable parameters, it is then defined as:
@@ IV-D Interpretability of Complex-Valued Kolmogorov-Arnold Networks
One major advantage of KANs is their inherent interpretability. Since the KAN learns on the edges, the learned univariate functions are directly applied to each input feature of the current layer. As proposed in [1, 2], we can utilize this property in two ways to explain the inner workings of the KAN: We can give importance scores to all edges and vertices to understand which parts of the network significantly influence the output and we can visualize the learned functions on the edges to understand their input-output relation. Both of these can also be utilized for CVKANs.
To calculate importance scores for edges and vertices, KAN 2.0 [2] uses the standard deviation of the output of that edge or vertex over the whole dataset distribution to annotate every edge and node with a relevance score by iterating backwards through the network. We adopt this idea and propose to use the standard deviation of the complex distribution. For and its mean, it is defined as:
@@ V Experiments
Just like in [1], we show that our network is capable of learning simple symbolic correlations between input and output based on synthetic formulae and outperforms its real-valued counterparts on complex-valued problems.
@@
For our experiments based on synthetic formulae we sample data points for each variable within our grid range and calculate the target output based on the symbolic formula. Our grid range is always in or in respectively. Because of Batch Normalization the resulting distribution has a standard deviation of and mean . We use a grid centered around zero spanning two standard deviations in each direction to optimize the trade off between a smaller grid size and minimizing the chance of the output of the former layer exceeding the grid. For the tabular knot dataset [40] we normalize each feature to be inside of our grid. If the dataset contains real-valued features we also feed them into our CVKAN as complex-valued numbers but set the imaginary part to zero.
@@ V-C Knot Classification
The knot dataset [40] originates from knot theory. Davies et al. calculated 15 invariants on different knots, from which 13 are real-and two complex-valued. Thus we need a real-valued KAN with 17 input features or a CVKAN with 15 complex-valued input features, out of which 13 have their imaginary part set to zero. The task on this dataset is to classify the knots based on their invariants into 14 different classes representing their signature. Therefore the output layer’s width of both real- and complex-valued KANs is set to to produce a probability vector. Since these probabilities are real-valued we need to learn a function in the output layer as described in subsection IV-A. The features of the dataset are normalized to fit in our grid of fixed size for real-valued KANs and for CVKAN.
@@
Fig. 4 shows the first layer of a CVKAN trained on the knot dataset with the opacity of nodes and edges showcasing the calculated relevance scores. Like in [1] the meridinal, here as one single complex number, and longitudinal translation are the most relevant features. To also quantitatively evaluate the explainability, we conduct a small study, where we exclude less relevant features and retrain the model on only the more relevant features. When reducing the dataset with this method to 7 or 3 features, we still obtain good results, only reducing the accuracy by and respectively (cf. Table VI). In contrast, when we leave out these 7 or 3 most relevant features, the performance drops to . This shows that this method determines the relevance of features accurately.
@@ VI Discussion
Our tool for visualizing the proposed CVKAN enables the understanding of the calculations happening inside our CVKAN and can thus help with (re-) discovering mathematical correlations in the data like Liu et al. pointed out in their real-valued KAN [1, 2]. The calculated importance scores reflect the true relevances of the features, as is shown by our experiments of training only on the or most important features or by leaving them out in Table VI. While the plot of our CVKAN trained on synthetic formula in Fig. 3 may seem unintuitive at first glance, the learned functions that are plotted show an interesting insight. In the bottom layer there exist two roots in both functions for and . The root points are exactly rotated around ° when comparing those two functions. This is caused by shifts that cancel out:
@@ VII Conclusion
As future work, we aim to apply CVKAN to real world problems with complex-valued data, such as finding solutions to complex-valued PDEs by utilizing them in Kolmogorov-Arnold-Informed Neural Networks [4]. Additionally, the concept of CVKAN could be extended further to other arbitrary hypercomplex algebras such as quaternions [41, 42].

### v4

@@ Abstract
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. The final published version (M. Wolff, F. Eilers and X. Jiang, ”CVKAN: Complex-Valued Kolmogorov-Arnold Networks,” 2025 International Joint Conference on Neural Networks (IJCNN), Rome, Italy, 2025, pp. 1-9, doi: 10.1109/IJCNN64981.2025.11227425) is available at IEEE Xplore: https://ieeexplore.ieee.org/document/11227425
I Introduction
The recently published Kolmogorov-Arnold Network (KAN) [liu2024kan, liu2024kan20] has proven to be a successful new approach for problems especially in the field of symbolic function fitting [yu2024kan] and solving physical equations [wang2025kolmogorov]. Adding to its success is the intrinsic explainability of KANs because of learnable univariate functions on edges instead of linear weights and fixed nonlinear activation functions in classical Multilayer Perceptrons. These learned univariate functions hold all model weights and can easily be visualized, thus making it intuitively understandable how single layers impact the model output.
Another emerging field in machine learning research are Complex-Valued Neural Networks (CVNNs), which have shown great success in a multitude of applications as well as good theoretical properties in terms of overfitting and generalization [lee2022complex, bassey2021survey]. Especially for applications with complex-valued input types, these models have shown promising results.
@@ I Introduction
The original KANs used fitting of B-Splines to learn the edge functions, however this training is rather slow and can be unstable, thus [li2024kolmogorov] proposed to replace the originally used B-Splines with easier to learn Radial Basis Functions. While it would be possible to utilize complex-valued B-Splines [ahlberg1967complex, unser2000fractional, forster2006complex] to construct a complex-valued KAN, we propose to adopt the idea of RBFs to the complex domain, to benefit from their easier and faster training process. Additionally, we propose to replace the tedious dynamic grid extension used in KANs by Batch Normalization to prevent the model to fall outside the predefined grid.
@@ II-A Kolmogorov-Arnold Networks
After the recent introduction of KANs by Liu et al. [liu2024kan, liu2024kan20] there has been a great variety of attempts to apply the ideas of KANs to different fields like image processing [bodner2024convolutional], satellite image segmentation [cambrin2024kan], graph neural networks [kiamari2024gkan, bresson2024kagnns] and even transformers [yang2024kolmogorov]. Hou et al. [hou2024comprehensive] give a detailed overview of the different applications and extensions of KANs. Their explainability can be of great value in fields where machine learning approaches are strongly regulated like survival analysis in medicine or engineering [knottenbelt2024coxkan]. Multiple works have already benchmarked the performance of KANs against MLPs [yu2024kan, poeta2024benchmarking] and found KANs to be a more suitable alternative in some fields. Alter et al. [alter2024robustness] found that large-scale KANs are more robust against adversarial attacks as MLPs and thus form an interesting direction for further research in multiple fields.
In [wang2025kolmogorov] the authors explore different partial differential equation forms based on KAN instead of MLP for solving forward and inverse problems in computational physics. A systematical comparison demonstrates that the KAN approach significantly outperforms MLP regarding accuracy and convergence speed. Further successful applications of KAN can be found for operator learning in computational mechanics [Abueidda2025] and image classification [Jamali2024].
@@ II-B Complex-Valued Neural Networks
After early introduction of CVNNs [aizenberg1971about, hirose2003complex, nitta1997extension, yang1994complex] they have lately risen in popularity since the introduction of building blocks for deep learning architectures [trabelsi2018deep]. Since then a lot of theoretical contributions have been made [tan2022real, eilers2023building, zhang2021optical] to enable a multitude of applications [zhong2023real, chen2023spectral, liu2023pixelwise, xing2023phase, yakupouglu2024comparison].
The most closely related prior work in the complex domain are deep Complex Radial Basis Function Neural Networks [chen2008fully, soares2024deep]. However, in CRBFNs, the RBFs are applied to all inputs of a neuron (e.g. vertex of the computational graph) simultaneously, while in CVKANs the RBFs are applied on the edges of the computational graph to each value individually. Thus CRBFNs are architecturally more similar to classical MLPs with RBFs as activation functions, where we instead aim to adopt the KAN framework to the complex domain.
@@ III Real-Valued Kolmogorov-Arnold Networks
The Kolmogorov-Arnold representation theorem (1) [kolmogorov1958On] states that any continuous function can be represented by a superposition of univariate functions.
This formula can be used to construct a two layer neural network with inputs, one output and learnable functions . For each layer and each feature an individual univariate function has to be learned, so combination of features is only done by summing the outputs of all these functions. While theoretically possible, this theorem was long deemed to be of little use for machine learning [girosi1989representation], because the inner functions can be highly nonsmooth.
However, Liu et al. [liu2024kan] have generalized this formulation of only one hidden layer of size and one single output to arbitrary network depths and layer widths with to overcome the problem of nonsmooth inner functions. Let be the -th neuron in layer , then every node in layer is connected to every node in the following layer by an edge with and . On every edge there is one learnable univariate function and the values of all incoming edges into a single node of the next layer are summed. This way the multiplication with linear weights between layers in classical MLPs has been replaced by learnable univariate functions and the nonlinear but fixed activation functions in the nodes of MLPs have been replaced by plain summation. In the original KAN the functions are learned using B-Splines [liu2024kan, liu2024kan20].
@@ III-A Real-Valued Radial Basis Functions
The way B-Splines are used to learn the functions in KANs [liu2024kan, liu2024kan20] is computationally intensive and slow. Thus different improved methods for learning a function have been proposed like DeepOKAN [Abueidda2025] or FastKAN [li2024kolmogorov], in which the authors suggest to replace B-Splines with RBFs for faster computation. Equation (2) describes a RBF that is symmetrical and centered around .
@@ IV-A Complex-Valued Radial Basis Functions
Chen et al. [chen2008fully] have chosen to treat real and imaginary parts separately and construct a complex-valued RBF out of two real-valued RBFs which only depend on the real or imaginary part of the input each. On the other hand Soares et al. [soares2024deep] use a single real-valued RBF that depends on both the real and imaginary parts of the input jointly. While the latter apply these multivariate RBFs inside vertices, we propose to use univariate RBFs on edges instead to maintain the idea of KANs and achieve higher explainability. We define a as:
@@ IV-B Residual Activation Function
We propose to use SiLU (8) as this residual activation function, a complex-valued equivalent to the SiLU-function (7) used in [liu2024kan, liu2024kan20, li2024kolmogorov]:
@@ IV-C Complex-Valued Batch Normalization
For KANs it is necessary to fix the grid range. However, it is not guaranteed that the output of an intermediate layer (and thus the input of the next layer) stays within a grid that was fixed before training. To tackle this challenge [liu2024kan] proposed dynamic extension of the grid during training. This is however a time intensive process that might slow down the learning process. Instead we propose to normalize the output of a vertex after summation. We propose to use BatchNorm () [ioffe2015batch] for the Radial Basis Function based FastKAN [li2024kolmogorov]. To adapt this to the complex domain, we explore three different complex-valued Batch Normalization approaches.
First, we adapt BatchNorm [trabelsi2018deep], where the covariance matrix of the complex distribution is normalized and an output distribution is learned through a learnable covariance (e.g. a symmetric positive definite) matrix. With being the mean and covariance over the batch dimension and learnable parameters, it is then defined as:
@@ IV-D Interpretability of Complex-Valued Kolmogorov-Arnold Networks
One major advantage of KANs is their inherent interpretability. Since the KAN learns on the edges, the learned univariate functions are directly applied to each input feature of the current layer. As proposed in [liu2024kan, liu2024kan20], we can utilize this property in two ways to explain the inner workings of the KAN: We can give importance scores to all edges and vertices to understand which parts of the network significantly influence the output and we can visualize the learned functions on the edges to understand their input-output relation. Both of these can also be utilized for CVKANs.
To calculate importance scores for edges and vertices, KAN 2.0 [liu2024kan20] uses the standard deviation of the output of that edge or vertex over the whole dataset distribution to annotate every edge and node with a relevance score by iterating backwards through the network. We adopt this idea and propose to use the standard deviation of the complex distribution. For and its mean, it is defined as:
@@ V Experiments
Just like in [liu2024kan], we show that our network is capable of learning simple symbolic correlations between input and output based on synthetic formulae and outperforms its real-valued counterparts on complex-valued problems.
@@
For our experiments based on synthetic formulae we sample data points for each variable within our grid range and calculate the target output based on the symbolic formula. Our grid range is always in or in respectively. Because of Batch Normalization the resulting distribution has a standard deviation of and mean . We use a grid centered around zero spanning two standard deviations in each direction to optimize the trade off between a smaller grid size and minimizing the chance of the output of the former layer exceeding the grid. For the tabular knot dataset [davies2021advancing] we normalize each feature to be inside of our grid. If the dataset contains real-valued features we also feed them into our CVKAN as complex-valued numbers but set the imaginary part to zero.
@@ V-C Knot Classification
The knot dataset [davies2021advancing] originates from knot theory. Davies et al. calculated 15 invariants on different knots, from which 13 are real-and two complex-valued. Thus we need a real-valued KAN with 17 input features or a CVKAN with 15 complex-valued input features, out of which 13 have their imaginary part set to zero. The task on this dataset is to classify the knots based on their invariants into 14 different classes representing their signature. Therefore the output layer’s width of both real- and complex-valued KANs is set to to produce a probability vector. Since these probabilities are real-valued we need to learn a function in the output layer as described in subsection IV-A. The features of the dataset are normalized to fit in our grid of fixed size for real-valued KANs and for CVKAN.
@@
Fig. 4 shows the first layer of a CVKAN trained on the knot dataset with the opacity of nodes and edges showcasing the calculated relevance scores. Like in [liu2024kan] the meridinal, here as one single complex number, and longitudinal translation are the most relevant features. To also quantitatively evaluate the explainability, we conduct a small study, where we exclude less relevant features and retrain the model on only the more relevant features. When reducing the dataset with this method to 7 or 3 features, we still obtain good results, only reducing the accuracy by and respectively (cf. Table VI). In contrast, when we leave out these 7 or 3 most relevant features, the performance drops to . This shows that this method determines the relevance of features accurately.
@@ VI Discussion
Our tool for visualizing the proposed CVKAN enables the understanding of the calculations happening inside our CVKAN and can thus help with (re-) discovering mathematical correlations in the data like Liu et al. pointed out in their real-valued KAN [liu2024kan, liu2024kan20]. The calculated importance scores reflect the true relevances of the features, as is shown by our experiments of training only on the or most important features or by leaving them out in Table VI. While the plot of our CVKAN trained on synthetic formula in Fig. 3 may seem unintuitive at first glance, the learned functions that are plotted show an interesting insight. In the bottom layer there exist two roots in both functions for and . The root points are exactly rotated around ° when comparing those two functions. This is caused by shifts that cancel out:
@@ VII Conclusion
As future work, we aim to apply CVKAN to real world problems with complex-valued data, such as finding solutions to complex-valued PDEs by utilizing them in Kolmogorov-Arnold-Informed Neural Networks [wang2025kolmogorov]. Additionally, the concept of CVKAN could be extended further to other arbitrary hypercomplex algebras such as quaternions [parcollet2020survey, lopez2024towards].